OpenStudy (anonymous):

The problem is:Minimize C=2x+4y subject to x+y(greater than or equal to)3, x+2y(greater than or equal to)4, x(greater than or equal to)0, y(greater than or equal to)0....I will rewrite it in a comment

OpenStudy (anonymous):

The problem is: Minimize C=2x+y\[x+y \ge 3, x+2y \ge 4, x \ge 0, y \ge 0\]